OpenStudy (olaniyiinjersey):

Let z=xy 2 +xy 3 1a. Find a tangent vector at (2,1,4) that is parallel to the XZ plane (i.e. projection of vector onto XY plane is parallel to the X axis b Find an equation to the tangent line at (2,1,4) that is parallel to the XZ plane (i.e. projection of vector onto XY plane is parallel to the X axis c Find a tangent vector at (2,1,4) that is parallel to the YZ plane (i.e. projection of vector onto XY plane is parallel to the Y axis d Find an equation to the tangent line at (2,1,4) that is parallel to the XZ plane (i.e. projection of vector onto XY plane is parallel to

OpenStudy (barinly):

@Festinger can u help me

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (irishboy123):

For the first part, the normal vector will be \(\vec n = \nabla \varphi\) where \(\varphi\) is the surface in level form \(\varphi = xy^2 + xy^3 - z = const\) so \(\vec n = <y^2 + y^3, 2xy + 3xy^2, -1>\) which you can compute for \((2,1,4)\) There will be an infinite number of tangent directions vectors \(\vec t = <t_1, t_2, t_3 > \) at the point but we know 2 things. \(\vec t \cdot \vec n = 0\) and \(t_2 = 0\) as \(\vec t\) is parallel with x-z plane so solve \(\vec n \cdot <t_1, 0, t_3 > = 0\) at the point \((2,1,4)\) to find out how \(t_1\) and \(t_3\) are related repeat for other bits ___________________________ you can also calc the normal vector for surface \( \vec r = \left[ \begin{matrix} x \\ y \\ xy^2 + xy^3 \end{matrix} \right ] \) as cross product of tangent vectors: \(\vec n = \dfrac{\partial \vec r }{\partial x} \times \dfrac{\partial \vec r }{\partial y} \) at the point,.....